A manufacturer’s annual losses follow a distribution with density function

f(x) = ( 2.5(0.6)2.5 /x3.5 , : x > 0.6

( 0, : otherwise.


To lower its losses, the manufacturer purchases an insurance policy with an annual deductible of 2. What is the mean of the manufacturer’s annual losses not paid by the insurance policy?


1
Expert's answer
2021-09-21T02:34:59-0400

Let us represent the manufacturers annual losses by a random variable X. Then it’s pdf f(x) is given as:

"f(x)=\left\{\begin{matrix}

\frac{2.5(0.6)^{2.5}}{x^{3.5}}\;for \;x>0.6 & \\ 

0\; otherwise & 

\end{matrix}\right."

Let us represent losses not paid by a new random variable Y, then we can write Y in terms of X as:

"Y=\left\{\begin{matrix}

x \; for \;0.6 < x< 2 & \\ 

2 \; for\; x \geqslant 2 & 

\end{matrix}\right."

Therefore, the mean of the manufacturer’s annual losses not paid will be equal to expected (or mean) value of the random variable Y.

Now let us recall following proposition.

If X is a continuous random variable with probability density function f(x) on interval [a,b] and h(x) be a function of x, then the expected value of h(x) is

"E[h(x)] = \int^b_a h(x) \cdot f(x) dx"

In above proposition, the random variable Y is equivalent of h(x), hence the expected value of Y can be given as:

"E(Y) = \int^2_{0.6}x \cdot f(x) dx + \int^{\infty}_2 2 f(x)dx \\

= \int^2_{0.6} x \times \frac{2.5(0.6)^2.5}{x^{3.5}}dx + \int^{\infty }_2 \frac{2 \times 2.5(0.6)^{2.5}}{x^{3.5}}dx \\

= \int^2_{0.6} \frac{0.6971}{x^{2.5}}dx + \int^{\infty }_2 \frac{1.3943}{x^{3.5}}dx \\

= 0.6971 \int^2_{0.6} x^{-2.5} dx + 1.3943 \int^{\infty }_2 x^{-.35} dx \\

= 0.6971 [\frac{x^{-1.5}}{-1.5}]^2_{0.6} + 1.3943 [\frac{x^{-2.5}}{-2.5}]^{\infty }_2 \\

= -0.4647 [x^{-1.5}]^2_{0.6} + (-0.5577)[x^{-2.5}]^{\infty }_2 \\

= -0.4647 [2^{-1.5} -0.6^{-1.5}] -0.5577 [0-2^{-2.5}] \\

= -0.4647 \times (-1.7981) -0.5577 \times (-0.1768) \\

= 0.93"